OpenStudy (anonymous):

Hi everyone! Since an equation is exact if dM/dy = dN/dy, then how do I show if -f(x)dx+g(y)dy=0 is exact? Not sure how to show this in general terms. Thanks! :o)

OpenStudy (anonymous):

I meant dM/dy = dN/dx

ganeshie8 (ganeshie8):

is f(x) a fnction of x alone ?

ganeshie8 (ganeshie8):

if so, the derivative with respect to `y` would be 0, yes ?

OpenStudy (anonymous):

uhm...well...the deriv of x with respect to y...yes it would be zero I suppose

ganeshie8 (ganeshie8):

similarly the derivative of g(y) with respect to `x` would be 0 since they both are equal to 0, they are equal to each other.

OpenStudy (anonymous):

interesting!...I didn't think that f(x) meant that "x" is actually the function...thinking of it this way, then yes they would both be zero! Thanks Ganeishi! :o)

OpenStudy (anonymous):

I have such a hard time spelling your name! X0/

ganeshie8 (ganeshie8):

I'm not so sure if I understood your question correctly hmm

ganeshie8 (ganeshie8):

can you provide the details of full problem??

OpenStudy (anonymous):

Well the whole question is simply prove that -f(x)dx+g(y)dy=0 is an exact equation

ganeshie8 (ganeshie8):

Ohkay.. then f(x) is a function of "x" alone no y terms in f(x)

ganeshie8 (ganeshie8):

M = -f(x) N = g(y) dM/dy = d/dy (-f(x)) = 0 dN/dx = d/dy (g(y)) = 0 that means dM/dy = dN/dx = 0 so the given equation is exact.

OpenStudy (anonymous):

got it! can you reason me through one more?

ganeshie8 (ganeshie8):

wil try, ask..

OpenStudy (anonymous):

i need to confirm whether or not y=tan(x) is a solution to dy/dx=1+y^2 on interval -pi/2,pi/2

OpenStudy (anonymous):

when I solve for y, i get y=arctan(x+c)

OpenStudy (anonymous):

is that correct?

OpenStudy (anonymous):

supposedly y=tan(x) is a solution on that interval but I can't seem to prove it

OpenStudy (anonymous):

:o(

OpenStudy (freckles):

I think the confirmation is you finding y' given y=tan(x) and seeing if the equation dy/dx=1+y^2 holds on the given interval

OpenStudy (freckles):

if y=tan(x) then y'=?

OpenStudy (anonymous):

uhm...nuts....isn't that sec^2(x) ?

OpenStudy (freckles):

\[y'=1+y^2\\ \sec^2(x)=1+\tan^2(x) \text{ is this equation true ? } \]

OpenStudy (anonymous):

lemme think a moment

OpenStudy (freckles):

hint: think Py...rean

OpenStudy (anonymous):

yes it's true

OpenStudy (freckles):

I purposely let out some letters there

OpenStudy (freckles):

Ok also about the equation you found it should have been arctan(y)=x+c where in this case they have c as 0 arctan(y)=x so then y=tan(x) which is what we wanted to show was a solution

OpenStudy (freckles):

But when I think confirmation I think about the first way I talked about just showing it is a solution instead of finding the solution

OpenStudy (anonymous):

wait...one second...lemme say something...one sec

OpenStudy (anonymous):

darn it...I just drew out the problem and it disappeared...one sec...sorry :o(

OpenStudy (freckles):

openstudy has been doing this thing to me where when I type things the page goes up while I'm typing so I can't see what I'm typing and that is annoying I don't know if that is why you accidentally deleted it or not but some times I wish openstudy was a person so it can hear me yell at it

OpenStudy (anonymous):

sorry...my pc kicked me off...lemme try again

OpenStudy (anonymous):

okay, it's working now...one sec...thanks for being patient with me! :o)

ganeshie8 (ganeshie8):

ok waiting

OpenStudy (anonymous):